Jay~Z is currently riding his new car moving at constant speed of 60 kph. If his destination is 200 km away from him, how fast he must accelerate (in m/s^2) to arrive their in exactly 10 minutes


1
Expert's answer
2021-08-17T10:26:05-0400

10 minutes is 600 seconds. 60 km/h is


"v=60·1000/3600=16.67\text{ m/s}."


200 km is 200 000 m.

To cover this distance in 600 s starting at 16.67 m/s, he must accelerate at


"d=v_0t+\frac{at^2}{2},\\\space\\
a=\frac{2(d-v_0t)}{t^2}=1.056\text{ m/s}^2."